OpenStudy (anonymous):

Determine whether or not the indicated set of 3 x 3 matrices is a subspace of M33. The set of all non singular 3 x 3 matrices?

OpenStudy (kirbykirby):

I think you can find an easy counter example for this. A non-singular matrix means it's invertible, so it also means that its determinant \(\ne 0\). Consider: \[ A=\left[\begin{matrix}1 & 0 & 0 \\ 0 & -1 & 0 \\ 0 & 0 & 1\end{matrix}\right]\] \[ B=\left[\begin{matrix}1 & 0 & 0 \\ 0 & 1 & 0 \\ 0 & 0 & 1\end{matrix}\right]\] Then, \[A+B =\left[\begin{matrix}1 & 0 & 0 \\ 0 & 0 & 0 \\ 0 & 0 & 1\end{matrix}\right]\] Notice that \(\det(A)=-1\), \(\det(B)=1\) but \(\det(A+B)=0\) so \(A+B\) is not closed under addition

OpenStudy (anonymous):

Hey thank you very much so I'm guessing if it is the case where it is a singular matrix means it is not invertible. So if I was asked the same question but for singular matrix's. So... \[A = \left[\begin{matrix}1 & 0 & 1\\ 0 & 1 & 0 \\ 0 & 0 & 0\end{matrix}\right]\] \[B = \left[\begin{matrix}0 & 0 & 0 \\ 0 & 0 & 0\\ 0 & 0 & 1\end{matrix}\right]\] Then \[A + B = \left[\begin{matrix}1 & 0 & 1\\ 0 & 1&0\\0&0&1\end{matrix}\right]\] and \[\det(A) = 0 , \det(B) = 0 , \det(A+B)=1\] so it is not closed under addition? I guess the big question I have is could I show this with \[cA\]?

OpenStudy (kirbykirby):

Yeah your example works :) It's not closed under addition, so the set of singular matrices wouldn't be a subspace either. You don't actually have to show the case for \(cA\) (closed under scalar multiplication). Just showing either not closed under addition, or not closed under scalar multiplication is enough to show it's not a subspace. (In fact, it's possible that it may be closed for scalar mult. and not closed for addition, or vice-versa, and it's still not a subspace... both operations have to be closed)

OpenStudy (anonymous):

Thank you so much